# MISTRAL: a model for radiatively efficient AGN winds in cosmological simulations

### v1

@@ Abstract
Feedback from active galactic nuclei (AGN) is crucial for regulating galaxy evolution. Motivated by observations of broad absorption line winds from rapidly accreting supermassive black holes (SMBHs), we introduce the mistral AGN feedback model, implemented in the arepo code. mistral comes in two versions: continuous radial (mistral-continuous) and stochastic bipolar momentum deposition (mistral-stochastic). Using the framework of the IllustrisTNG simulations, we explore the effect of mistral on BH and galaxy properties, through an idealized Milky Way-mass galaxy and cosmological zoom simulations run down to $z=2$. Unlike standard thermal AGN feedback prescriptions, mistral generates galaxy-scale winds that mimic outflows driven by BH accretion. mistral-continuous produces short-lived galactic fountains, and is inefficient at regulating the growth of massive galaxies at $z=2$. In contrast, mistral-stochastic efficiently suppresses star formation in massive galaxies, and reproduces the empirical stellar-to-halo mass and ($z=0$) BH-stellar mass relations. By supporting large-scale ($>50\,\rm kpc$) outflows while simultaneously preventing gas inflows, mistral-stochastic additionally regulates the cold and hot gas fractions at both galaxy and halo scales. mistral-stochastic therefore works self-consistently across the halo mass range explored $\left(10^{12}-3\times 10^{13}\,\rm M_{\odot}\right)$, without adopting a SMBH-mass dependent AGN feedback scheme such as the one used in IllustrisTNG. Our model is a promising tool for predicting the impact of radiatively efficient AGN winds on galaxy evolution, and interpreting the growing population of high-redshift galaxies and quasars observed by JWST. This work is part of the "Learning the Universe" collaboration, which aims to infer the physical processes governing the evolution of the Universe.
@@
pubyear: 2025pagerange: MISTRAL: a model for radiatively efficient AGN winds in cosmological simulationsB
@@ 1 Introduction
For more than two decades, most (if not all) massive galaxies are thought to host a supermassive black hole (SMBH) in their centres, with masses ranging from a million to a few billion solar masses (Magorrian et al., 1998; Schödel et al., 2002). As material accretes onto SMBHs, vast amounts of energy are released in the form of energetic radiation, winds and jets (Fabian, 2012; Tombesi et al., 2013; Cicone et al., 2014; Genzel et al., 2014), powering so-called active galactic nuclei (AGN). Since the energy released by a SMBH can easily exceed the binding energy of its host galaxy (e.g. King & Pounds, 2015), AGN are thought to play a significant role in galaxy evolution (Silk & Rees, 1998; Kauffmann & Haehnelt, 2000; Croton et al., 2006; Somerville et al., 2008; Somerville & Davé, 2015; Naab & Ostriker, 2017). Depending on the efficiency with which the BH accretion energy released couples to the surrounding gas, the back reaction of an AGN – referred to as AGN feedback – can affect both BH growth and the evolution of the host galaxy by heating and ejecting gas, and (or) preventing gaseous inflows, and thus depriving both star formation and BH accretion of their source of fuel (Fabian, 2012; Gaspari et al., 2012; Harrison, 2017). This BH-galaxy coevolution scenario is further supported by the tight correlations observed between the BH mass and the properties of their host galaxies, such as stellar bulge luminosity, mass and stellar velocity dispersion (Tremaine et al., 2002; Kormendy & Ho, 2013; Reines & Volonteri, 2015, even if also non-causal origins of these relations have been discussed by e.g. Hirschmann et al., 2010; Jahnke & Macciò, 2011).
@@
Conversely, in high-luminosity AGN (with bolometric luminosities of $10^{46}-10^{48}\,\rm erg\,s^{-1}$), the accreting mass is efficiently converted into radiation, driving radiatively efficient winds. These winds are powered through radiation, line or magnetic pressure, or interactions with free electrons or dust. They can originate on small scales within the SMBH torus or accretion disk, in the broad line region, or at larger nuclear scales. AGN winds are observed via broad absorption lines (BAL) in the rest-frame UV, as well as broad optical, UV and infrared emission lines in the spectra of quasars (e.g. King & Pounds, 2015), which are characteristic of strong outflows of molecular and ionised gas (e.g. Richards et al., 2011; Harrison et al., 2014; Cicone et al., 2014; Hamann et al., 2018; Wylezalek et al., 2020; Herrera-Camus et al., 2020). In particular, BAL winds, found in $\sim 20-50$ percent of the quasar population across all redshifts (e.g Hewett & Foltz, 2003; Dai et al., 2008; Bischetti et al., 2023; Maiolino et al., 2024), are considered a key driver of AGN feedback; they are often highly energetic, with outflow velocities ranging from a few thousand to a few tens of thousands of $\rm km\,s^{-1}$ (e.g. Fiore et al., 2017; Choi et al., 2020; Rodríguez Hidalgo et al., 2020), and can lead to outflows that extend from parsec to galactic scales, where the latter can reach several kpc from the central SMBH or even further (e.g. Arav et al., 2018; Choi et al., 2022).
By transferring mass, momentum and energy from the nuclear region to the interstellar medium (ISM) and the circumgalactic medium (CGM) of galaxies, radiatively efficient winds expel and heat large amounts of gas, and therefore have the ability to impact star formation (e.g. Veilleux et al., 2020, for a review). This AGN feedback mode is thought to be particularly relevant at cosmic noon ($z\simeq 2$), when both cosmic star formation density and quasar activity peak (Madau & Dickinson, 2014). This scenario is further supported by the prevalence of AGN-driven circum-nuclear outflows observed at this epoch (e.g. Förster Schreiber et al., 2014; Genzel et al., 2014), indicating the importance of radiatively efficient AGN feedback (also referred to as quasar-mode feedback in the literature) in the early assembly of galaxies.
@@
To smooth over the small-scale complexity, the common approach to modelling AGN feedback in cosmological simulations is to rely on subgrid prescriptions, meant to reproduce the consequences of processes that happen at unresolved scales. Initially, AGN feedback was implemented by simply assuming that a fraction of the rest-mass SMBH accreted energy couples thermally to the surrounding gas (Springel et al., 2005; Di Matteo et al., 2005), which was shown to be insufficient in reducing cooling and star formation in most massive systems with stellar masses beyond $10^{11}\rm\,M_{\odot}$ (at the typical resolution achieved in cosmological simulations). To overcome this, more complex prescriptions have been progressively developed (Choi et al., 2017; Bourne & Sijacki, 2017; Talbot et al., 2021; Huško et al., 2022; Rennehan et al., 2024), though some of these are still prohibitively expensive to include in large cosmological simulations. In such coarse resolution simulations, isotropic thermal BH energy deposition remains a widely adopted approach to implementing feedback in the radiatively efficient BH regime (e.g. Di Matteo et al., 2012; Vogelsberger et al., 2014; Hirschmann et al., 2014; Dubois et al., 2014; Springel et al., 2018, but see Schaye et al., 2015; Choi et al., 2015; Davé et al., 2019).
To better regulate gas fractions and star formation in the most massive galaxies, some simulations transition to another AGN feedback mode when the Eddington ratio drops below a certain threshold ($\sim 10^{-2}$), injecting kinetic energy (Dubois et al., 2012; Weinberger et al., 2017), or thermal energy with a higher efficiency (Di Matteo et al., 2012; Steinborn et al., 2015) or in the form of AGN-inflated bubbles of hot gas (Sijacki et al., 2007). Other simulations, which do not distinguish between radiatively efficient and inefficient regimes (Schaye et al., 2015, 2023), distribute the thermal AGN energy from a reservoir of feedback energy by stochastically heating gas elements to a high temperature (Booth & Schaye, 2009). To reproduce galaxy observables at $z=0$, cosmological simulations often enhance the effectiveness of their AGN thermal mode (Schaye et al., 2015), or implement a two-mode feedback scheme with an accretion rate-based switch (Dubois et al., 2012; Weinberger et al., 2017; Davé et al., 2019). The latter approach is adopted for instance in the Illustris-TNG simulations (TNG hereafter, Springel et al., 2018; Pillepich et al., 2018a; Naiman et al., 2018; Nelson et al., 2018; Marinacci et al., 2018; Nelson et al., 2019a).
@@
In addition, it remains unclear whether radiatively efficient AGN feedback in the form of thermal energy deposition can generate realistic galactic-scale gaseous outflows, as observed out to $z=7.5$ (Liu et al., 2024, and references therein). Thus, motivated by observations of BAL winds, Ostriker et al. (2010) and Choi et al. (2012, 2015) first implemented a momentum/kinetic energy based AGN feedback model for generating winds driven by radiatively efficient BH accretion. Using similarly motivated prescriptions (Anglés-Alcázar et al., 2017; Davé et al., 2019; Costa et al., 2020), recent simulations suggest that quasar-driven outflows can play an important role in galaxies. Notably, they contribute to efficiently regulating star formation and X-ray luminosities (Choi et al., 2018; Costa et al., 2018; Davé et al., 2019), bringing the $z=0$ galaxy stellar mass function and cold and hot gas fractions in simulations closer to observations, compared to incorporating only thermal quasar energy injection (Sijacki et al., 2007; Vogelsberger et al., 2013).
@@
Motivated by this success, it is now within reach to investigate how radiatively efficient AGN winds affect galaxy and BH evolution out to early cosmic epochs. In this paper, we build on the AGN feedback model of Ostriker et al. (2010) and Choi et al. (2012) and introduce the mistral [^1] Mistral is the name of a cold and strong wind that blows in South of France. subgrid model for cosmological simulations, implemented into the arepo code (Springel, 2010). mistral bridges the gap between radiatively efficient AGN wind prescriptions in high-resolution, idealised simulations (e.g. Costa et al., 2020; Sivasankaran et al., 2024), and traditional subgrid models implemented in large volume cosmological simulations. As such, mistral allows us to investigate the role of radiatively efficient AGN-driven winds in galaxy evolution. In particular, in this paper, we use mistral to explore to what extent AGN-driven winds affect BH and their host galaxy properties around cosmic noon.
@@
For this purpose, in this study, we couple mistral with the well-tested galaxy formation physics from the TNG cosmological simulations. For reproducing realistic galaxy populations at $z=0$, the TNG simulations distinguish between a thermal isotropic AGN energy input at high Eddington ratios (referred to as the TNG thermal model), and a kinetic prescription in the low-accretion regime (introduced by Weinberger et al., 2017). Our aim with mistral is to incorporate a physically motivated prescription for radiatively efficient AGN winds, in order to understand their impact on the BH-galaxy co-evolution compared to previous, standard AGN feedback models. This is also one of the key goals of the "Learning the Universe" collaboration [^2] https://learning-the-universe.org, of which this work is a part. With this intention in mind, we have explored two variants of our model implementation: a continuous radial deposition of momentum (mistral-continuous), and stochastic bipolar momentum injection (mistral-stochastic) – the arepo analogue of the model implemented by Choi et al. (2012) in the gadget code (Springel, 2005). In this study, both versions are used independently, and at all Eddington ratios.
@@ 2.1.1 Black hole accretion
where $m_{i}/m_{\rm target}$ is the mass of the gas cell $i$ compared to the target gas mass resolution enforced by the refinement scheme of the arepo code. $w_{\rm k}(r_{i})$ is the cubic spline smoothing kernel weight associated with this gas cell $i$, which scales with its distance $r_{i}$ to the BH. Essentially, this means that the BH smoothing length may vary over time, in order to enclose a similar number of cells $n_{\rm BH,ngb}\pm 4$. In TNG50, TNG100 and TNG300 (the TNG simulations with box side lengths of $\sim 50$, $100$ and $300\,\rm cMpc$ respectively), $n_{\rm BH,ngb}$ is set to 512, 256 and 128 (see also Appendix B in Weinberger et al., 2017). In what follows, we refer to the sphere of radius $h_{\rm BH}$ as the BH smoothing volume. In this work, the value of $h_{\rm BH}$ varies from a few hundred $\rm pc$ to a few $\rm kpc$. This parameter increases when the gas density near the BH decreases (as a result of the gas cell width expanding), which preferentially occurs after efficient AGN feedback episodes.
@@ 2.1.2 AGN feedback in the radiatively efficient regime
Particularly relevant in the high-accretion regime, when the bolometric luminosity is the highest, Vogelsberger et al. (2013) introduced a radiative, electro-magnetic AGN feedback channel that is included in the TNG simulations. This phenomenological model aims at capturing the effect of the ionising radiation emitted by SMBHs on the thermal state of the gas. For this purpose, the fraction of AGN bolometric luminosity that does not couple to the gas is combined with a redshift-dependent UV background (Faucher-Giguère et al., 2009), which is then used to infer metal line cooling and heating rates from pre-tabulated tables. To determine their net cooling rate, each gas cell is assigned a total bolometric intensity, derived from the obscured bolometric luminosity of all the BHs from the simulation, depending on their distance to the gas cell. This radiative feedback channel renders the TNG thermal feedback more efficient, as star formation is slightly more suppressed in the high-accretion regime compared to when no AGN radiative feedback is included (see Fig. 15 from Vogelsberger et al., 2013).

### v2

# MISTRAL: a model for AGN winds from radiatively efficient accretion in cosmological simulations

@@ Abstract
Feedback from active galactic nuclei (AGN) is crucial for regulating galaxy evolution. Motivated by observations of broad absorption line winds from rapidly accreting supermassive black holes (SMBHs), we introduce the mistral AGN feedback model, implemented in the arepo code. mistral comes in two versions: continuous radial (mistral-continuous) and stochastic bipolar momentum deposition (mistral-stochastic). Using the framework of the IllustrisTNG simulations, we explore the effect of mistral on BH and galaxy properties, through an idealized Milky Way-mass galaxy and cosmological zoom simulations run down to $z=2$. Unlike standard thermal AGN feedback prescriptions, mistral generates galaxy-scale winds that mimic outflows driven by BH accretion. mistral-continuous produces short-lived galactic fountains, and is inefficient at regulating the growth of massive galaxies at $z=2$. In contrast, mistral-stochastic efficiently suppresses star formation in massive galaxies, reproduces the empirical stellar-to-halo mass relation, and yields a consistent trend of BH-stellar mass evolution. By supporting large-scale outflows while simultaneously preventing gas inflows, mistral-stochastic additionally regulates the cold and hot gas fractions at both galaxy and halo scales. mistral-stochastic therefore works self-consistently across the halo mass range explored $\left(10^{12}-3\times 10^{13}\,\rm M_{\odot}\right)$, without adopting a SMBH-mass dependent AGN feedback scheme such as the one used in IllustrisTNG. Our model is a promising tool for predicting the impact of AGN winds on galaxy evolution, and interpreting the growing population of high-redshift galaxies and quasars observed by JWST. This work is part of the "Learning the Universe" collaboration, which aims to infer the physical processes governing the evolution of the Universe.
@@
pubyear: 2025pagerange: MISTRAL: a model for AGN winds from radiatively efficient accretion in cosmological simulationsB
@@ 1 Introduction
For more than two decades, most (if not all) massive galaxies are thought to host a supermassive black hole (SMBH) in their centres, with masses ranging from a million to a few billion solar masses (Magorrian et al., 1998; Schödel et al., 2002). As material accretes onto SMBHs, vast amounts of energy are released in the form of energetic radiation, winds and jets (Fabian, 2012; Tombesi et al., 2013; Cicone et al., 2014; Genzel et al., 2014), powering so-called active galactic nuclei (AGN). Since the energy released by a SMBH can easily exceed the binding energy of its host galaxy (e.g. King & Pounds, 2015), AGN are thought to play a significant role in galaxy evolution (Silk & Rees, 1998; Kauffmann & Haehnelt, 2000; Bower et al., 2006; Croton et al., 2006; Somerville et al., 2008; Somerville & Davé, 2015; Naab & Ostriker, 2017). Depending on the efficiency with which the BH accretion energy released couples to the surrounding gas, the back reaction of an AGN – referred to as AGN feedback – can affect both BH growth and the evolution of the host galaxy by heating and ejecting gas, and (or) preventing gaseous inflows, and thus depriving both star formation and BH accretion of their source of fuel (Fabian, 2012; Gaspari et al., 2012; Harrison, 2017). This BH-galaxy coevolution scenario is further supported by the tight correlations observed between the BH mass and the properties of their host galaxies, such as stellar bulge luminosity, mass and stellar velocity dispersion (Tremaine et al., 2002; Kormendy & Ho, 2013; Reines & Volonteri, 2015, even if also non-causal origins of these relations have been discussed by e.g. Hirschmann et al., 2010; Jahnke & Macciò, 2011).
@@
Conversely, in high-luminosity AGN (with bolometric luminosities of $10^{46}-10^{48}\,\rm erg\,s^{-1}$), the accreting mass is efficiently converted into radiation, driving winds from radiatively efficient accretion disks. These winds are powered through magnetic or radiation pressure on lines, free electrons or dust. They can originate on small scales within the SMBH torus or accretion disk, in the broad line region, or at larger nuclear scales. AGN winds are observed via broad absorption lines (BAL) in the rest-frame UV, as well as broad optical, UV and infrared emission lines in the spectra of quasars (e.g. King & Pounds, 2015), which are characteristic of strong outflows of molecular and ionised gas (e.g. Richards et al., 2011; Harrison et al., 2014; Cicone et al., 2014; Hamann et al., 2018; Wylezalek et al., 2020; Herrera-Camus et al., 2020). In particular, BAL winds, found in $\sim 20-50$ percent of the quasar population across all redshifts (e.g Hewett & Foltz, 2003; Dai et al., 2008; Bischetti et al., 2023; Maiolino et al., 2024), are considered a key driver of AGN feedback; they are often highly energetic, with outflow velocities ranging from a few thousand to a few tens of thousands of $\rm km\,s^{-1}$ (e.g. Fiore et al., 2017; Choi et al., 2020; Rodríguez Hidalgo et al., 2020), and can lead to outflows that extend from parsec to galactic scales, where the latter can reach several kpc from the central SMBH or even further (e.g. Arav et al., 2018; Choi et al., 2022).
By transferring mass, momentum and energy from the nuclear region to the interstellar medium (ISM) and the circumgalactic medium (CGM) of galaxies, winds from radiatively efficient accretion expel and heat large amounts of gas, and therefore have the ability to impact star formation (e.g. Veilleux et al., 2020, for a review). This AGN feedback mode is thought to be particularly relevant at cosmic noon ($z\simeq 2$), when both cosmic star formation density and quasar activity peak (Madau & Dickinson, 2014). This scenario is further supported by the prevalence of AGN-driven circum-nuclear outflows observed at this epoch (e.g. Förster Schreiber et al., 2014; Genzel et al., 2014), indicating the importance of radiatively efficient AGN feedback (also referred to as quasar-mode feedback in the literature) in the early assembly of galaxies.
@@
To smooth over the small-scale complexity, the common approach to modelling AGN feedback in cosmological simulations is to rely on subgrid prescriptions, meant to reproduce the consequences of processes that happen at unresolved scales. Initially, AGN feedback was implemented by simply assuming that a fraction of the rest-mass SMBH accreted energy couples thermally to the surrounding gas (Springel et al., 2005; Di Matteo et al., 2005), which was shown to be insufficient in reducing cooling and star formation in most massive systems with stellar masses beyond $10^{11}\rm\,M_{\odot}$ (at the typical resolution achieved in cosmological simulations). To overcome this, more complex prescriptions have been progressively developed (Choi et al., 2017; Bourne & Sijacki, 2017; Talbot et al., 2021; Huško et al., 2022; Rennehan et al., 2024; Huško et al., 2024, 2025), though some of these are still prohibitively expensive to include in large cosmological simulations. In such coarse resolution simulations, isotropic thermal BH energy deposition remains a widely adopted approach to implementing feedback in the radiatively efficient BH regime (e.g. Di Matteo et al., 2012; Vogelsberger et al., 2014; Hirschmann et al., 2014; Dubois et al., 2014; Springel et al., 2018, but see Schaye et al., 2015; Choi et al., 2015; Davé et al., 2019).
To better regulate gas fractions and star formation in the most massive galaxies, some simulations transition to another AGN feedback mode when the Eddington ratio drops below a certain threshold ($\sim 10^{-2}$), injecting kinetic energy (Dubois et al., 2012; Weinberger et al., 2017), or thermal energy with a higher efficiency (Di Matteo et al., 2012; Steinborn et al., 2015) or in the form of AGN-inflated bubbles of hot gas (Sijacki et al., 2007). Other simulations, which do not distinguish between radiatively efficient and inefficient regimes (Schaye et al., 2015, 2023), distribute the thermal AGN energy from a reservoir of feedback energy by stochastically heating gas elements to a high temperature (Booth & Schaye, 2009). To reproduce galaxy observables at $z=0$, cosmological simulations often enhance the effectiveness of their AGN thermal mode, for instance by stochastically heating a selected number of cells to a large enough temperature to overcome excessive radiative losses (Schaye et al., 2015), or implement a two-mode feedback scheme with an accretion rate-based switch (Dubois et al., 2012; Weinberger et al., 2017; Davé et al., 2019). The latter approach is adopted for instance in the Illustris-TNG simulations (TNG hereafter, Springel et al., 2018; Pillepich et al., 2018a; Naiman et al., 2018; Nelson et al., 2018; Marinacci et al., 2018; Nelson et al., 2019a).
@@
In addition, it remains unclear whether radiatively efficient AGN feedback in the form of thermal energy deposition can generate realistic galactic-scale gaseous outflows, as observed out to $z=7.5$ (Liu et al., 2024, and references therein). Thus, motivated by observations of BAL winds, Ostriker et al. (2010) and Choi et al. (2012, 2015) first implemented a momentum/kinetic energy based AGN feedback model for generating winds driven by radiatively efficient BH accretion. Using similarly motivated prescriptions (Anglés-Alcázar et al., 2017; Davé et al., 2019; Costa et al., 2020), recent simulations suggest that quasar-driven outflows can play an important role in galaxies. Notably, they contribute to efficiently regulating star formation and X-ray luminosities (Choi et al., 2018; Costa et al., 2018; Davé et al., 2019), bringing the $z=0$ galaxy stellar mass function and cold and hot gas fractions in simulations closer to observations, compared to incorporating only continuous thermal quasar energy injection (Springel et al., 2005; Sijacki et al., 2007).
@@
Motivated by this success, it is now within reach to investigate how AGN winds from radiatively efficient accretion affect galaxy and BH evolution out to early cosmic epochs. In this paper, we build on the AGN feedback model of Ostriker et al. (2010) and Choi et al. (2012) and introduce the mistral [^1] Mistral is the name of a cold and strong wind that blows in South of France. subgrid model for cosmological simulations, implemented into the arepo code (Springel, 2010). mistral bridges the gap between radiatively efficient AGN wind prescriptions in high-resolution, idealised simulations (e.g. Costa et al., 2020; Sivasankaran et al., 2025), and traditional subgrid models implemented in large volume cosmological simulations. As such, mistral allows us to investigate the role of radiatively efficient AGN-driven winds in galaxy evolution. In particular, in this paper, we use mistral to explore to what extent AGN-driven winds affect BH and their host galaxy properties around cosmic noon.
@@
For this purpose, in this study, we couple mistral with the well-tested galaxy formation physics from the TNG cosmological simulations. For reproducing realistic galaxy populations at $z=0$, the TNG simulations distinguish between a thermal isotropic AGN energy input at high Eddington ratios (referred to as the TNG thermal model), and a kinetic prescription in the low-accretion regime (introduced by Weinberger et al., 2017). Our aim with mistral is to incorporate a physically motivated prescription for AGN winds from radiatively efficient accretion, in order to understand their impact on the BH-galaxy co-evolution compared to previous, standard AGN feedback models. This is also one of the key goals of the "Learning the Universe" collaboration [^2] https://learning-the-universe.org, of which this work is a part. With this intention in mind, we have explored two variants of our model implementation: a continuous radial deposition of momentum (mistral-continuous), and stochastic bipolar momentum injection (mistral-stochastic) – the arepo analogue of the model implemented by Choi et al. (2012) in the gadget code (Springel, 2005). In this study, both versions are used independently, and at all Eddington ratios.
@@ 2.1.1 Black hole accretion
where $m_{i}/m_{\rm target}$ is the mass of the gas cell $i$ compared to the target gas mass resolution enforced by the refinement scheme of the arepo code. $w_{\rm k}(r_{i})$ is the cubic spline smoothing kernel weight associated with this gas cell $i$ (in units of inverse volume), which scales with its distance $r_{i}$ to the BH. Essentially, this means that the BH smoothing length may vary over time, in order to enclose a similar number of cells $n_{\rm BH,ngb}\pm 4$. In TNG50, TNG100 and TNG300 (the TNG simulations with box side lengths of $\sim 50$, $100$ and $300\,\rm cMpc$ respectively), $n_{\rm BH,ngb}$ is set to 512, 256 and 128 (see also Appendix B in Weinberger et al., 2017). In what follows, we refer to the sphere of radius $h_{\rm BH}$ as the BH smoothing volume. In this work, the value of $h_{\rm BH}$ varies from a few hundred $\rm pc$ to a few $\rm kpc$. This parameter increases when the gas density near the BH decreases (as a result of the gas cell width expanding), which preferentially occurs after efficient AGN feedback episodes.
@@ 2.1.2 AGN feedback in the radiatively efficient regime
Particularly relevant in the high-accretion regime, when the bolometric luminosity is the highest, Vogelsberger et al. (2013) introduced what they refer to as a radiative, electro-magnetic AGN feedback channel. This channel consists of depositing thermal energy, and is included in the TNG simulations. More specifically, this phenomenological model aims at capturing the effect of the ionising radiation emitted by SMBHs on the thermal state of the gas. For this purpose, the fraction of AGN bolometric luminosity that does not couple to the gas is combined with a redshift-dependent UV background (Faucher-Giguère et al., 2009), which is then used to infer metal line cooling and heating rates from pre-tabulated tables. To determine their net cooling rate, each gas cell is assigned a total bolometric intensity, derived from the obscured bolometric luminosity of all the BHs from the simulation, depending on their distance to the gas cell. This radiative feedback channel renders the TNG thermal feedback more efficient, as star formation is slightly more suppressed in the high-accretion regime compared to when no AGN radiative feedback is included (see Fig. 15 from Vogelsberger et al., 2013).